OpenStudy (anonymous):

. Describe the type of solutions you can get when you solve an absolute value equation. Give an example of each type

OpenStudy (anonymous):

You can get a positive or negative solution. For instance |x + 6| = 3 x + 6 = -3 x = -9 Or x + 6 = 3 x = -3 Here we have two negatives for our solution.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Here is one with a positive and negative solution |2x - 1| = 7 2x - 1 = 7 2x = 7+1 2x = 8 2x/2 = 8/2 x = 4 OR |2x - 1| = -7 2x - 1 = -7 2x = -7+1 2x = -6 2x/2 = -6/2 x = -3 Here we have a positive 4 and a negative 3 .... {4, -3}